Two angles are complementry. The sum of the measures of the first angle and half of the second angle is 89.5 degrees. Find the measurements of each angle. Please show me the solution.
Answer by checkley77(12844)

x+(90-x)/2=89.5
(2x+90-x)/2=89.5
x+90=2*89.5
x+90=179
x=179-90
x=89 degrees.
90-89=1 degree.
Proof:
89+(90-89)/2=89.5
89+1/2=89.5
89.5=89.5